Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2024 logbook, Contact: F.Wolff-Fabris, A.Galler, H.Sinn  Not logged in ELOG logo
Message ID: 104     Entry time: 13 May 2024, 10:36
Author: Romain Letrun 
Group: SPB 
Subject: Shift summary 

X-ray delivery

  • 11.56 keV, 1-197 pulses @ 564 kHz, 2.4-3.1 mJ.
  • Beam down a bit over 1 hour due to accelerator interlock problem.

Optical laser delivery

  • 840 nm, 50 fs, 50 pulses @ 141 kHz.
  • Stable delivery.

Achievements/Observations

  • PP laser timed.
  • Completed TR-SFX dataset at 30 ps delay.

Issues

  • Scantool stuck in ACQUIRING state. Recovered by restarting the device.
  • HIREX data missing for most of the night. GOTTHARD II control and receiver devices found down in the morning. Server on which they are reunning appears to have crashed at 1:36. DOC informed in the morning.
ELOG V3.1.4-7c3fd00